A fertilizer is to be diluted with water so that the fertilizer is just 1% of the solution. How much water should be mixed with 5cm^3 of fertilizer?
Answer by ptaylor(2198)


Let x=amount of water needed
Now we know that the amount of pure fertilizer before the mixture takes place (5 cm^3)(Note: 5cm^3 *100% or 5cm^3*1 or 5cm^3) has to equal the amount of pure fertilizer in the final mixture 0.01(5+x) so our equation to solve is:
5=0.01(5+x) get rid of parens
5=0.05+0.01x subtract 0.05 from each side
5-0.05=0.05-0.05+0.01x collect like terms
4.95=0.01x divide each side by 0.01
x=495 cm^3------------------------amount of water needed

CK
5cm^3=0.01(5cm^3+495cm^3)
5cm^3=0.01(500cm^3)
5cm^3=5cm^3
Hope this helps-----ptaylor